DEUTERATED METHANOL IN LOW-MASS PROTOSTARS

\underline{BERENGERE PARISE}, Centre d’Etude Spatiale des Rayonnements, 9 avenue du Colonel Roche, BP 4346, 31028 Toulouse Cedex 4, France ; CECILIA CECCARELLI, LAOG, BP 53, 38041 Grenoble Cedex 09, France; A.G.G.M. TIELENS, SRON, PO Box 800, 9700 AV Groningen, The Netherlands; ERIC HERBST, Department of Physics, The Ohio State University, 174 W. 18th Ave. Columbus, OH 43210-1106, USA; BERTRAND LEFLOCH, LAOG, BP 53, 38041 Grenoble Cedex 09, France; EMMANUEL CAUX, CESR, 9 avenue du Colonel Roche, BP 4346, 31028 Toulouse Cedex 4, France ; ALAIN CASTETS, Observatoire de Bordeaux, BP 89, 33270 Floirac, France; INDRA MUKHOPAD-HYAY, College of Natural Sciences, Dakota State University, Madison, SD 57042-1799, USA; LAU-RENT PAGANI, LERMA & FRE 2460 du CNRS, Observatoire de Paris, 61 Av. de l’Observatoire de Paris, 75014 Paris, France; LAURENT LOINARD, Instituto de Astronomia, UNAM, Apdo Postal 72-3 (Xangari), 58089 Morelia, Michoacan, Mexico.

In the last five years, deuterated species have been observed to be very abundant in low-mass protostars envelopes. Recent laboratory studies of deuterated isotopomers of methanol provided astronomers with the frequencies and line strengths, allowing us to observe these species in astronomical sources. Doubly-deuterated methanol was first detected towards the low-mass protostar IRAS16293-2422, with a fractionation CHD_{2}OH/CH_{3}OH of 0.2. Subsequent observations showed that methanol is similarly deuterated in other low-mass protostars. These observations bring a very constraining tool to understand the processes leading to molecular deuteration in protostellar envelopes. Ongoing observations of other deuterated isotopomers of methanol will also be discussed in this talk.